OpenStudy (anonymous):

Solve the equation using factoring. x2 + 8x - 20 = 0 Anwsers: {-2, -10} {2, 10} {-2, 10} {2, -10

OpenStudy (anonymous):

Find the two numbers that multiply to -20, and add to +8.

OpenStudy (anonymous):

i think i got it...-2, 10?

OpenStudy (anonymous):

But then that represents (x-2)(x+10) = x^2 + 8x - 20, so we aren't done yet.

OpenStudy (anonymous):

aww. no more medals for wrong answers?

OpenStudy (anonymous):

Then we need to solve x-2 = 0, and x+10 = 0

OpenStudy (anonymous):

yes.. they try lol

OpenStudy (anonymous):

I can assure you is one of those ;D keep thinking :p

OpenStudy (anonymous):

@dpaInc LOL , I would vote for the quadratic equation formula, it always works! \[\Large x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] in your case you have to substitute ... a=1 b=8 c=-20 calculate , and there's nothing else to get confused O_O

OpenStudy (anonymous):

that is definitely right...